a ball is thrown horizontally from the roof of a building 46.0 m tall and lands 25.1 m from the base. what was its initial speed?
1
Expert's answer
2012-01-31T09:48:56-0500
Let's find the time of flight T:

H = g*T²/2 ==> T = (2*H/g)^(1/2) = (2*46/9.8)^(1/2) ≈ 3.0639 s.

So, the initial speed was

V = L/T = 25.1/3.0639 ≈ 8.1920 m/s.